OpenStudy (anonymous):

3x-y-3z if x=-2, y=1, and z=-2

OpenStudy (kenrhi):

Are you just wanting an answer? 3(-2) - 1 - 3(-2) = -6-1+6=-1.....-1

OpenStudy (anonymous):

\[(3)(−2)−1−(3)(−2)= -1]/ it is = -1